OpenStudy (anonymous):

Consider the equation: X^3 - (1 + cosө + sinө) x^2 – ( cosөsinө + cosө + sinө)x – cosөsinө = 0, whose roots are a,b,c. 1. Find a^2 + b^2 + c^2. 2. Number of values of ө є [0,2π] such that the equation has at least two equal roots. 3. The greatest possible difference between the roots.

OpenStudy (campbell_st):

well looking at your problem knowing the roots your polynomial P(x) = (x -a)(x-b)(x -c) if you expand distribute to get you polynomial you will be able to compare the coefficients of each term given \[P(x) = x^3 -(1 \cos(\theta) + \sin(\theta)x^2 -(\cos(\theta)\sin(\theta)+\cos(\theta)+\sin(\theta))x - \cos(\theta)\sin(\theta)\] and looking at your 1st problem \[(a + b + c)(a + b + c) = a^2 + b^2 + c^2 + 2ab + 2ac + 2bc\] so you need to manipulate the equation so that you are looking at \[(a + b + c)^2 - something\] hope this makes sense